# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v1

@@ Abstract
Keywords: Bayesian analysis; Gaussian graphical models; Finite mixture of regression models; Heterogeneous network estimation.
@@ 1 Introduction
For cancer, heterogeneity is a fundamental characteristic and is essential for unraveling the inherent complexities of biological systems. To describe and accommodate heterogeneity, researchers have conducted extensive studies on subgroup identification with high-dimensional molecular data, which can be roughly classified into unsupervised and supervised. Unsupervised approaches typically emphasize differences in the means of molecular factors (Figure 1a), with examples including [17] and [4], and variances (Figure 1b), as exemplified by [24]. Under the supervised analysis framework, a popular approach is the finite mixture of regression models, which captures the heterogeneity of genetic effects on clinical outcomes (Figure 1c). Examples include both Bayesian approaches [19, 1] and penalization approaches [12, 22, 28]. These models enable the integration of information from both clinical outcomes and molecular data to identify clinically relevant subgroups. Despite significant achievements, existing techniques remain inadequate as they focus on the distributional properties or the influences of individual variables but neglect the interconnections among variables. There is a strong need for more informative approaches for subgroup identification.
Figure 1: A comparative overview of unsupervised and supervised approaches for subgroup identification.
@@
The importance of network-based subgroup identification has been emphasized in recent literature [20]. Cancer-specific networks are pivotal in understanding complex biological processes [2]. They utilize information embedded in the interconnections among variables (Figure 1d), providing richer insights than marginal means and variances. As illustrated in the literature [6, 25], in addition to heterogeneity, there can be commonalities across subgroup-specific networks, which can help improve estimation efficiency and enhance interpretability through information borrowing.
@@
Among the diverse statistical methods for network analysis, Gaussian graphical models (GGMs) have been prominent. Here, the nodes of a network are represented by random variables with a joint Gaussian distribution, while the edges reflect conditional dependencies between nodes that are determined by non-zero entries of the inverse covariance matrix (or precision matrix) [16, 9]. To detect unknown heterogeneity, several network-based methods have been proposed under the framework of the Gaussian mixture model (GMM), with particular attention to identifying both structural differences and potential commonalities among multiple networks. For example, [10] incorporated a joint fused graphical lasso penalty into GMM for learning the related structures across different unknown subgroups. Similarly, a SCAN method proposed by [11] used a joint group graphical lasso penalty for simultaneous subgroup identification and joint graphical model estimation. [18] decomposed the GGM into a set of sparse regression problems and introduced an approach for jointly estimating multiple networks in the presence of unknown sample heterogeneity. Under the Bayesian framework, [5] proposed a BCGS method that imposes a joint spike-and-slab graphical lasso prior on the precision matrices when the subgroup information is unknown.
@@
The remainder of the paper is structured as follows. In Section 2, we present the proposed Bayesian model formulation and prior specification and the EM algorithm for posterior inference. Theoretical results for the proposed estimator are also established in this section. In Section 3, we investigate the performance of the proposed approach using simulated data. In Section 4, we apply the proposed approach to real-world data. Concluding remarks are offered in Section 5. The proofs for the results stated in Section 2 are in the online Supplementary Materials.
@@ 2 Methods
We consider data with a survival outcome. Modeling survival presents more challenges due to non-negativity and censoring, and similar approaches can be applied to other clinical outcomes. Let $T$, $\Delta$, and $\bm{X}=\left(X_{1},\ldots,X_{p}\right)^{\top}$ represent the logarithm of observed survival time, censoring indicator, and $p$-dimensional vector of predictors (e.g., gene expressions), respectively. Here, $T=\operatorname{min}\left(Z,C\right)$ with $Z$ and $C$ being the logarithms of survival and censoring times, and $\Delta={I}\{Z\leq C\}$ with ${I}\{\cdot\}$ being the indicator function. Assume there are $n$ subjects $\left\{\left(t_{i},\delta_{i},\bm{x}_{i}\right),i=1,\ldots,n\right\}$, which form $K$ subgroups. Denote $\bm{z},\bm{c},\bm{\delta},$ and $\bm{t}$ as the vectors of $z_{i}$’s, $c_{i}$’s, $\delta_{i}$’s, and $t_{i}$’s, respectively, and X as the $n\times p$ matrix composed of $\bm{x}_{i}$’s.
@@ 2.1 Bayesian mixture modeling
In (1), to accommodate censoring, we adopt prior $z_{i}^{\delta_{i}}c_{i}^{1-\delta_{i}}$ for $t_{i}$. Here, we assume random censoring with $c_{i}$ being independent of $z_{i}$ and $\bm{x}_{i}$, as in the literature [22, 8]. We propose combining the AFT model with the mixture Gaussian graphical model to accommodate the multi-layer heterogeneity. Specifically, in the AFT model, we introduce a subgroup-specific coefficient $\bm{\beta}_{k}$ to explicitly capture the heterogeneous associations of the molecular predictors with survival time. Here, a reparameterization technique based on the noise level is adopted for $\bm{\beta}_{k}$ to ensure that the proposed estimator remains scale-invariant under the mixture model [23, 28]. This technique facilitates an “adaptive adjustment” of the coefficients, aligning them with data characteristics and leading to more accurate estimations. On the other hand, within the Gaussian graphical model, we utilize $\bm{\mu}_{k}$ and $\bm{\Omega}_{k}$ to characterize heterogeneity attributed to both differential mean expressions and differential network structures. Consequently, the network’s heterogeneity is not only intricately linked to the interested measurements but also closely connected to the distribution of survival time, making the resulting networks clinically more meaningful. For example, researchers can better understand the differences between patients with good survival and those with bad survival on specific biological networks, thereby more effectively guiding subsequent treatments.
@@ 2.2 Prior specification
$\omega_{k,jl}\mid\gamma_{k,jl}\sim\gamma_{k,jl}\operatorname{LP}\left(\omega_{k,jl};v_{1}\right)+\left(1-\gamma_{k,jl}\right)\operatorname{LP}\left(\omega_{k,jl};v_{0}\right),\gamma_{k,jl}\sim\operatorname{Bern}\left(p_{1}\right),\text{ for }j\neq l,$
@@
Here, $\gamma_{k,jl}$ is a binary indicator corresponding to $\omega_{k,jl}$, $\operatorname{LP}(y;v)=\exp(-|y|/v)/(2v)$ is the density function of the Laplace distribution with a scale parameter $v$, $p_{1}\in(0,1)$ is the prior probability for the Bernoulli distribution, and $\bm{\pi}_{g}=\left(\pi_{1},\ldots,\pi_{K}\right)^{\top}$ is the vector composed of the Multinomial distribution’s parameters, which specifies the prior probabilities that each sample is drawn from different subgroups. $\lambda_{1}$, $\lambda_{2}$, and $v_{1}>v_{0}$ are the positive tuning parameters.
In the proposed Bayesian framework, we first introduce a spike-and-slab Laplace prior [21, 26] for $\omega_{k,jl}$ based on the indicator $\gamma_{k,jl}$, for identifying individual sparse patterns of the $K$ networks. With this prior, when $\gamma_{k,jl}=1$, the corresponding entry $\omega_{k,jl}$ comes from the slab component with a relative large variance $2v_{1}^{2}$ and is likely to take a value away from zero; when $\gamma_{k,jl}=0$, $\omega_{k,jl}$ comes from the spike component with a small variance $2v_{0}^{2}$ that can induce values towards zero. Compared to the Laplace prior, the spike-and-slab Laplace prior involves two different scale parameters $v_{1}$ and $v_{0}$ to distinguish active from ignorable coefficients and can nonlinearly adapt to coefficients with $\gamma_{k,jl}$, achieving “selective shrinkage” with a smaller bias. This Laplace distribution-based prior can automatically achieve sparsity via the $L_{1}$ norm on $\omega_{k,jl}$. This advances from the commonly adopted spike-and-slab Gaussian prior which yields nonsparse posterior modes for $\omega_{k,jl}$’s and relies on a post-data thresholding-based selection strategy utilizing $\gamma_{k,jl}$. The strategy that is based on the estimation and sparsity of the continuous $\omega_{k,jl}$ can lead to more effective identification of sparse posterior patterns [21] and can also facilitate a simpler theoretical study and a more efficient computation.
@@ 2.3 Additional priors for accommodating common structures
$\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},u\left(\bm{L}^{(jl)}\right)^{-1}\right),\bm{\theta}_{jl}\mid\bm{\omega}_{jl}\sim{I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\},$
with $u$ being the positive tuning parameter. Here, we define $L^{(jl)}_{k,k}=(K-1)/\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)$ and $L^{(jl)}_{k,k^{\prime}}=-1/\sqrt{\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)\left(\theta_{k^{\prime},jl}^{2}+\epsilon^{2}\right)}$ for $1\leq k\neq k^{\prime}\leq K$ and $1\leq j<l\leq p$, with $\epsilon$ being a small constant. Thus, for each $\bm{\theta}_{jl}$, the prior involves $\bm{\theta}_{jl}^{\top}\bm{L}^{(jl)}\bm{\theta}_{jl}=0.5\sum_{k\neq k^{\prime}}\left(\theta_{k,jl}/\sqrt{\theta_{k,jl}^{2}+\epsilon^{2}}-\theta_{k^{\prime},jl}/\sqrt{\theta_{k^{\prime},jl}^{2}+\epsilon^{2}}\right)^{2}$, which is an approximation to the Bayesian counterpart of the sign-based constraint penalty $0.5\sum_{k\neq k^{\prime}}\left[\operatorname{sgn}(\theta_{k,jl})-\operatorname{sgn}(\theta_{k^{\prime},jl})\right]^{2}$. Combined with the indicator function ${I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\}$, the proposed approach encourages the pairwise relationships of predictors across subgroups to have similar signs, leading to networks with a shared pattern of sparsity as well as edges with shared positive or negative relationships. This significantly advances from the previous research based on the group penalty [6, 11, 14, 5], which utilizes group selection without appropriately accounting for the similarity of the estimated values and fails to sufficiently address the relationships between different networks [15]. This is also more flexible than the strategy based on fused lasso [6, 25, 10, 27], which promotes both shared directions and magnitudes of edges. Similar prior assumptions can also be applied to $\bm{\beta}$ and $\bm{\mu}$ when considering the common structures within regression coefficients and predictors’ means.
@@ 2.4 Posterior Inference
×$\times f\left(\bm{\gamma}\right)f\left(\bm{\theta}\mid\bm{\Omega}\right)f\left(\bm{0}\mid\bm{\theta}\right)f\left(\bm{g}\mid\bm{\pi_{g}}\right).$

### v2

@@ Abstract
Keywords: Bayesian analysis; Finite mixture of regression models; Gaussian graphical model; Heterogeneous network estimation.
@@ 1 Introduction
For cancer, heterogeneity is a fundamental characteristic and is essential for unraveling the inherent complexities of biological systems. To describe and accommodate heterogeneity, researchers have conducted extensive studies on subgroup identification with high-dimensional molecular data, which can be roughly classified into unsupervised and supervised. Unsupervised approaches typically emphasize differences in the means of molecular factors (Table 1, a), with examples including [18] and [4], and variances (Table 1, b), as exemplified by [26]. Under the supervised analysis framework, a popular approach is the finite mixture of regression models, which captures the heterogeneity of genetic effects on clinical outcomes (Table 1, d). Examples include both Bayesian approaches [21, 1] and penalization approaches [24, 31]. These models enable the integration of information from both clinical outcomes and molecular data to identify clinically relevant subgroups. Despite significant achievements, existing techniques remain inadequate as they focus on the distributional properties or the influences of individual variables but neglect the interconnections among variables. There is a strong need for more informative approaches for subgroup identification.
Table 1: Comparison of unsupervised and supervised subgroup identification models.
Type Model and Example
Unsupervised a. Mean-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{V})$, where $\bm{\mu}_{k}$ is the mean vector in the $k$th subgroup and $\bm{V}$ is the diagonal covariance matrix. Example: $\bm{\mu}_{1}=(1,0,0)^{\top},\ \bm{\mu}_{2}=(0,1,-1)^{\top}$.
b. Variance-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{D}_{k})$, where $\bm{D}_{k}$ is the diagonal covariance matrix in the $k$th subgroup. Example: $\bm{D}_{1}=\text{diag}(1,2,1),\ \bm{D}_{2}=\text{diag}(2,1,2)$.
c. Network-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{\Omega}_{k}^{-1})$, where $\bm{\Omega}_{k}$ is the precision matrix in the $k$th subgroup. Example: $\bm{\Omega}_{1}=\begin{pmatrix}1&-1&0\\ -1&2&0\\ 0&0&1\end{pmatrix},\ \bm{\Omega}_{2}=\begin{pmatrix}2&0&-1\\ 0&1&0\\ -1&0&2\end{pmatrix}$.
Supervised d. Linear effect-based: $Z\sim\mathcal{N}(\bm{\beta}_{k}^{\top}\bm{X},\sigma_{k}^{2})$, where $\bm{\beta}_{k}$ and $\sigma_{k}$ are the regression parameter vector and standard deviation in the $k$th subgroup. Example: $\bm{\beta}_{1}=(-1,0,0)^{\top},\ \bm{\beta}_{2}=(0,1,-1)^{\top};\ \sigma_{1}=1,\ \sigma_{2}=2$.
@@
The importance of network-based subgroup identification has been emphasized in recent literature [22]. Subtype-specific networks are pivotal in understanding complex biological processes. They utilize information embedded in the interconnections among variables (Table 1, c), providing richer insights than marginal means and variances. As illustrated in the literature [7, 19], in addition to heterogeneity, there can be commonalities across subgroup-specific networks, which can help improve estimation efficiency and enhance interpretability through information borrowing.
@@
Among the diverse statistical methods for network analysis, Gaussian graphical models (GGMs) have been prominent. Here, the nodes of a network are represented by random variables with a joint Gaussian distribution, while the edges reflect conditional dependencies between nodes that are determined by non-zero entries of the inverse covariance matrix (or precision matrix) [16, 10, 5]. To detect unknown heterogeneity, several network-based methods have been proposed under the framework of the Gaussian mixture model (GMM), with particular attention to identifying both structural differences and potential commonalities among multiple networks. For example, [11] incorporated a joint fused graphical lasso penalty into GMM for learning the related structures across different unknown subgroups. Similarly, a SCAN method proposed by [12] used a joint group graphical lasso penalty for simultaneous subgroup identification and joint graphical model estimation. [20] proposed decomposing the GGM into a set of sparse regression problems and introduced a composite minimax concave penalty for jointly estimating multiple networks. Under the Bayesian framework, [6] proposed a BCGS method that imposes a joint spike-and-slab graphical lasso prior on the precision matrices.
@@ 2 Methods
We consider data with a survival outcome. Modeling survival presents more challenges due to non-negativity and censoring, and similar approaches can be applied to other clinical outcomes. Let $T$, $\Delta$, and $\bm{X}=\left(X_{1},\ldots,X_{p}\right)^{\top}$ represent the logarithm of the minimum of survival time and censoring time, censoring indicator, and $p$-dimensional vector of predictors (e.g., gene expressions), respectively. Here, $T=\operatorname{min}\left(Z,C\right)$ with $Z$ and $C$ being the logarithms of survival time and censoring time, and $\Delta={I}\{Z\leq C\}$ with ${I}\{\cdot\}$ being the indicator function. Assume there are $n$ subjects $\left\{\left(t_{i},\delta_{i},\bm{x}_{i}\right),i=1,\ldots,n\right\}$, which form $K$ subgroups. Denote $\bm{z},\bm{c},\bm{\delta},$ and $\bm{t}$ as the vectors of $z_{i}$’s, $c_{i}$’s, $\delta_{i}$’s, and $t_{i}$’s, respectively, and X as the $n\times p$ matrix composed of $\bm{x}_{i}$’s.
@@ 2.1 Bayesian mixture modeling
In (1), we assume random censoring with $c_{i}$ being independent of $z_{i}$ and $\bm{x}_{i}$, as in the literature [24]. We propose combining the AFT model with the mixture Gaussian graphical model to accommodate the multi-layer heterogeneity. Specifically, in the AFT model, we introduce a subgroup-specific coefficient $\bm{\beta}_{k}$ to explicitly capture the heterogeneous associations of the molecular predictors with survival time. Here, a reparameterization technique based on the noise level is adopted for $\bm{\beta}_{k}$ to ensure that the proposed estimator remains scale-invariant under the mixture model [25, 31]. This technique facilitates an “adaptive adjustment” of the coefficients, aligning them with data characteristics and leading to more accurate estimations. On the other hand, within the Gaussian graphical model, we utilize $\bm{\mu}_{k}$ and $\bm{\Omega}_{k}$ to characterize heterogeneity attributed to both differential mean expressions and differential network structures. Consequently, the network’s heterogeneity is not only intricately linked to the interested measurements but also closely connected to the distribution of survival time, making the resulting networks clinically more meaningful. For example, researchers can better understand the differences between patients with good survival and those with bad survival on specific biological networks, thereby more effectively guiding subsequent treatments. Here, to simplify our approach, consistent with prior studies [12, 8], we assume that the subgroup number $K$ is identifiable.
@@ 2.2 Prior specification
${\omega_{k,jj}\sim\operatorname{Exp}(\omega_{k,jj};\tau_{0})},\text{ for }j=1,\ldots,p,$
$\omega_{k,jl}\mid\gamma_{k,jl}\sim\gamma_{k,jl}\operatorname{LP}\left(\omega_{k,jl};v_{1}\right)+\left(1-\gamma_{k,jl}\right)\operatorname{LP}\left(\omega_{k,jl};v_{0}\right),\gamma_{k,jl}\sim\operatorname{Bern}\left(p_{1}\right),\text{ for }j<l,$
@@
Here, $\operatorname{Exp}(\omega_{k,jj};\tau_{0})$ is the Exponential distribution with parameter $\tau_{0}$, $\gamma_{k,jl}$ is a binary indicator corresponding to $\omega_{k,jl}$, $\operatorname{LP}(y;v)$ is the Laplace distribution with a scale parameter $v$, $p_{1}\in(0,1)$ is the prior probability for the Bernoulli distribution, and $\bm{\pi}_{g}=\left(\pi_{1},\ldots,\pi_{K}\right)^{\top}$ is the vector composed of the Multinomial distribution’s parameters, which specifies the prior probabilities that each sample is drawn from different subgroups. $\tau_{0}$, $\lambda_{1}$, $\lambda_{2}$, and $v_{1}>v_{0}$ are the positive tuning parameters.
In the proposed Bayesian framework, we first introduce a weakly informative Exponential prior to the diagonal elements $\omega_{k,jj}$’s of the precision matrices, which are not subject to shrinkage, following [10]. We introduce a spike-and-slab Laplace prior [23, 28] for $\omega_{k,jl}$ based on the indicator $\gamma_{k,jl}$, for identifying individual sparse patterns of the $K$ networks. With this prior, when $\gamma_{k,jl}=1$, the corresponding entry $\omega_{k,jl}$ comes from the slab component with a relative large variance $2v_{1}^{2}$ and is likely to take a value away from zero; when $\gamma_{k,jl}=0$, $\omega_{k,jl}$ comes from the spike component with a small variance $2v_{0}^{2}$ that can induce values towards zero. Compared to the Laplace prior, the spike-and-slab Laplace prior involves two different scale parameters $v_{1}$ and $v_{0}$ to distinguish active from ignorable coefficients and can nonlinearly adapt to coefficients with $\gamma_{k,jl}$, achieving “selective shrinkage” with a smaller bias. This Laplace distribution-based prior can automatically achieve sparsity via the $L_{1}$ norm on $\omega_{k,jl}$. This advances from the commonly adopted spike-and-slab Gaussian prior which yields nonsparse posterior modes for $\omega_{k,jl}$’s and relies on a post-data thresholding-based selection strategy utilizing $\gamma_{k,jl}$. The strategy that is based on the estimation and sparsity of the continuous $\omega_{k,jl}$ can lead to more effective identification of sparse posterior patterns [23] and can also facilitate a simpler theoretical study and a more efficient computation.
@@ 2.3 Additional priors for accommodating common structures
$\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},\left(u\bm{L}^{(jl)}\right)^{-1}\right),\bm{\theta}_{jl}\mid\bm{\omega}_{jl}\sim{I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\},$
with $u$ being the positive tuning parameter. Here, $\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},\left(u\bm{L}^{(jl)}\right)^{-1}\right)$ can be interpreted as a generative model in which a zero vector is observed from a Gaussian distribution parameterized by $\bm{\theta}_{jl}$, i.e., $f\left(\bm{0}\mid\bm{\theta}_{jl}\right)=\sqrt{u\left|\bm{L}^{(jl)}\right|/(2\pi)^{K}}\exp\left\{-u\left(\bm{0}-\bm{\theta}_{jl}\right)^{\top}\bm{L}^{(jl)}\left(\bm{0}-\bm{\theta}_{jl}\right)/2\right\}$. This interpretation follows the hierarchical framework of [30], where priors are constructed through pseudo-observations. We define $L^{(jl)}_{k,k}=(K-1)/\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)$ and $L^{(jl)}_{k,k^{\prime}}=-1/\sqrt{\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)\left(\theta_{k^{\prime},jl}^{2}+\epsilon^{2}\right)}$ for $1\leq k\neq k^{\prime}\leq K$ and $1\leq j<l\leq p$, with $\epsilon$ being a small constant. Thus, for each $\bm{\theta}_{jl}=(\theta_{1,jl},\ldots,\theta_{K,jl})^{\top}$, the prior involves $\bm{\theta}_{jl}^{\top}\bm{L}^{(jl)}\bm{\theta}_{jl}=0.5\sum_{k\neq k^{\prime}}\left(\theta_{k,jl}/\sqrt{\theta_{k,jl}^{2}+\epsilon^{2}}-\theta_{k^{\prime},jl}/\sqrt{\theta_{k^{\prime},jl}^{2}+\epsilon^{2}}\right)^{2}$, which is an approximation to the Bayesian counterpart of the sign-based constraint penalty, that is $0.5\sum_{k\neq k^{\prime}}\left[\operatorname{sgn}(\theta_{k,jl})-\operatorname{sgn}(\theta_{k^{\prime},jl})\right]^{2}$. Combined with the indicator function ${I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\}$, the proposed approach encourages the pairwise relationships of predictors across subgroups to have similar signs, leading to networks with a shared pattern of sparsity as well as edges with shared positive or negative relationships. This significantly advances from the previous research based on the group penalty [7, 12, 14, 6], which utilizes group selection without appropriately accounting for the similarity of the estimated values and fails to sufficiently address the relationships between different networks [15]. This is also more flexible than the strategy based on fused lasso [7, 11, 29], which promotes both shared directions and magnitudes of edges. Similar prior assumptions can also be applied to $\bm{\beta}$ and $\bm{\mu}$ when considering the common structures within regression coefficients and predictors’ means.
@@ 2.4 Posterior Inference
×$\times f\left(\bm{\gamma}\right)f\left(\bm{\theta}\mid\bm{\Omega}\right)f\left(\bm{0}\mid\bm{\theta}\right)f\left(\bm{g}\mid\bm{\pi_{g}}\right),$